OpenStudy (anonymous):

please help me solve (1/3)(7a-1)<2a+7

OpenStudy (johnweldon1993):

\[\large (\frac{1}{3})(7a - 1) < 2a + 7\] If we multiply both sides of this equation by 3...we can get rid of that (1/3) in front \[\large \cancel{3 \times \frac{1}{3}}(7a - 1) < 3(2a + 7)\] Now leets distribute the 3 into those parenthesis \[\large 7a - 1 < 6a + 21\] Now subtract 6a from both sides of ths equation...and also add 1 to both sides \[\large a < 22\]